OpenStudy (anonymous):

Complete the following ordered pairs for the equation. 4x + 5y = -22

OpenStudy (anonymous):

the equation represent a line since it is of first degree... in slope-intercept form... \[4x+5y=-22\]\[5y=-4x-22\]\[y=\frac{-4x-22}{5}\]\[y=-\frac{4}{5}x-\frac{22}{5}\]where slope \(m=-\frac{4}{5}\) and y-intercept \(-\frac{22}{5}\)

OpenStudy (anonymous):

plug in any values of x to solve equivalent values in y... that will complete ordered pairs....